OpenStudy (anonymous):

write expression as a single natural logarithm ln 3 - 5 ln 3

terenzreignz (terenzreignz):

Need to recall some of the properties of (natural) logarithms .... First... \[\large p\ln \ b = \ln \ b^p\]

terenzreignz (terenzreignz):

And second, \[\large \ln \ M - \ln \ N = \ln \ \frac{M}{N} \ \ \ \ \ \ \ N\ne0\]

OpenStudy (anonymous):

idk confuse

terenzreignz (terenzreignz):

Sorry, I didn't get notified of your response :) First, let's deal with the 5ln 3 part Using the first property I mentioned, this is equal to \[\large \ln \ 3^5\]right?

OpenStudy (anonymous):

yea so would it be like L3 - 5 LN 3 3x5x3=? or is that off

OpenStudy (anonymous):

Or no multiplication

terenzreignz (terenzreignz):

No multiplication, logarithm is something different. so let's start by gradually simplifying the expression \[\huge \ln \ 3 - 5\ln \ 3\] By the first property I gave you, this is equal to \[\huge \ln \ 3 - \ln \ 3^5\] right?

OpenStudy (anonymous):

yea I gotcha what would be next step instead of the multiplaction like the + problems would we divide

OpenStudy (anonymous):

or is that the final product

terenzreignz (terenzreignz):

Yes, you divide. Use the second property I gave you. (sorry if my replies are really late, I don't think I get notifications of your replies :( )